# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v5

@@ III-B Resolvability Analysis
$\begin{split}&\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{z},0})\\ =&\int_{z}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{X}_{sw})\\ &\times\log\frac{\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{X}_{sw})}{MKQ^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z},\end{split}$ (67)
@@
Note that the three terms in (III-B) are all random variables as they are functions of $\bm{X}_{sw},s=1,\cdots,K,w=1,\cdots,M$, which are randomly generated from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. Our strategy of bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is as follows. We directly investigate the expectation of each term in (III-B). By showing that the expectation of each term in (III-B) is small, we prove that the probability that a set of legitimate codewords can be obtained by the random coding scheme is positive. From the linearity of the expectation, we have
@@
Then, to upper bound the term (70), we divide the product space $\bm{X}\times\bm{Z}$ into two parts. For the first part, we show that the inner term $\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)$ for each $(\bm{X},\bm{Z})$ pair is under control. For the second part, we prove that the probability of it is exponentially decreasing. In fact, a precise upper bounding of the first part depends on the choice of the sizes of codebook and the key, which will be illustrated later in Theorem 6. Before establishing Lemma 9, some technical preparations are provided in Lemma 7 and Lemma 8. Define $\mathcal{E}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\leq\gamma_{1}\}$. Note that in the definition of $\mathcal{E}$, we use ${Q}^{(n)}_{\bm{Z},1}$ but not $\bar{Q}^{(n)}_{\bm{Z},1}$. For the ratio $\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}$, the following lemma is useful for the follow-up analysis, the proof of which can be found in Appendix D.
@@
For the event $\mathcal{E}^{c}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$, the following lemma provides an upper bound on its probability with proper chosen $\gamma_{1}$. Similar with Lemma 4, our strategy is separately considering the codewords with equal power and choosing an universal threshold $\gamma_{1}$. As a consequence, the sphere symmetry can be leveraged to show that the probability of the event $\{\bm{z}:\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$ for each $\bm{x}$ has large-deviation bounds.
@@
$\begin{split}&\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)\\ +&\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{z_{i}^{2}}{\sigma^{2}+\mu\Psi(n)}-\frac{(z_{i}-\sqrt{\Upsilon})^{2}}{\sigma^{2}}\right]\\ =&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\sum_{i=1}^{n}\frac{1}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\\ \times&\left[\mu\Psi(n)z_{i}^{2}-2(\sigma^{2}+\mu\Psi(n))\sqrt{\Upsilon}z_{i}+(\sigma^{2}+\mu\Psi(n))\Upsilon\right]\end{split}$ (78)
@@
The strategy of the proof of Lemma 8 is similar as Lemma 4. We first consider the codewords with the same $l^{2}$ norm $\sqrt{n\Upsilon}$ (on the sphere with the same radius or with equal power constraint), separately. For codewords on each sphere, we can utilize the sphere symmetry to simplify the analysis. Again, the parameter $\mu$ makes the interval under control so that an universal threshold $\gamma_{1}$ can be obtained. The inner probability $\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ for each $\bm{x}$ is a tail probability and can be bounded by concentration inequality from Chernoff technique. As the bounds of the tail probability depends only on the radius of the codewords and are all exponentially decreasing with $n$, we can get the maximum of the bounds for all $\bm{x}$, which is again exponential decreasing with $n$.
@@ III-C Existence of the Codebooks and Achievability Bound
By using McDiarmid’s Inequality with Lemma 10, we have the following lemma, the proof of which is omitted here.
@@
We have illustrated the random coding scheme, decoding method in Section III-A. The upper bound on the first type of decoding error in Lemma 4 and the average of the second type of decoding error has been analyzed in Lemma 6. We also bounded the expectation of the discriminative metric - $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. A legitimate covert code should be based on the previous lemma, the choice of various parameters both reliability and covertness will be illustrated in our main results as next theorem. We will provide characterizations on the size $M_{n}$ of each codebook for covert communication over AWGN channels with sufficient large $n$ through proper choice of the power level $\Psi(n)$. The theorem also characterizes the necessary number of key bits $K_{n}$, and it is shown that $K_{n}$ depends not only on the noise level $\sigma^{2}$ at the adversary’s channel and covert parameter $\delta$, but also depends on the hyper-parameter of truncation $\mu$.
@@
Under the above conditions, with sufficiently large $n$, we can choose $c_{0}=13$, $c_{1}=1$, $c_{2}=11$, $c_{3}=4$, and $c_{5}=3$ and a proper power level $c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}$. We obtain
$+\sigma(2\mu)^{1/2}(n\delta)^{1/4}$ $(\log e)^{3/4}\Phi^{-1}(\epsilon)+o(n^{\frac{1}{4}}),$ (114)
and the number of key bits should satisfy
@@
For each index $s$, we expurgate the $(1-\lambda_{2})M_{n}$ codewords with largest $\varepsilon_{sw}^{(2)}$ for from each sub-codebook $\{\bm{x}_{sw},w\in\langle M\rangle\}$. Thus,we get a new family of codebooks, each of which has size $\lambda_{2}M$. Without loss of generality, we assume that the remaining codewords have indexes from $1$ to $\lambda_{2}M$, and denote the maximum of error probability of the second type in these new codes as $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. These codewords with indexes from $\lambda_{2}M+1$ to $M$ have error probabilities of the second type at least larger than $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. As the average error probability of the second type for each old codebook is less than $\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. We have for each $s$, the following inequality holds,
@@
For the covertness condition in (108), from (113) and Lemma 3, the term $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ satisfies
@@
$\log$ $\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}$
@@
$\log M^{*}(n,\epsilon,\delta)\geq$ $\frac{n}{2}\log(1+\mu c_{4}\frac{\sqrt{n}}{n})+\frac{(\mu^{2}-\mu)c_{4}\frac{\sqrt{n}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})}\cdot\log e+\sqrt{n\cdot\frac{2\mu^{2}c_{4}\frac{\sqrt{n}}{n}+\frac{c_{4}^{2}\mu^{2}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})^{2}}}\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\log e$
$+$ $O(\log n)$
@@
With $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get a lower bound of $\log M^{*}(n,\epsilon,\sigma,\delta)$ in (III-C). The condition in (111) is equivalent to
@@
$+$ $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}$
$\times$ $\Phi^{-1}(1-\epsilon_{1})+c_{6}\log n\}$ (129)
@@
  • If $\sigma^{2}\leq\frac{1}{\mu^{2}}$, i.e., the channel between Alice and Willie is less noisy than the main channel, we have
@@
  • If $\sigma^{2}>\frac{1}{\mu^{2}}$, then $\phi(\mu,\sigma^{2})<0$ and the first order asymptotics of the optimal number of key bits should be $O(1)$.
@@
The support set of the codewords is $\{\bm{x}:\sqrt{\mu^{2}n\Psi(n)}\leq\|\bm{x}\|\leq\sqrt{n\Psi(n)}\}$. The upper bound of $\|\bm{x}\|$ is for the power constraint, and we set the inner bound $\sqrt{\mu^{2}n\Psi(n)}$ here for lower bounding the decoding error probability. If there exists codewords with too small $l^{2}$ norm, then there exists no $\gamma$ to lower bound the first type of error in decoding. In fact, the bounding of the first type of error in Lemma 4 relies on the fact that the power of each codeword lies in a finite interval (See Remark 5). Here we choose truncated Gaussian with decreasing variance $\mu\Psi(n)$. The parameter $\mu$ acts as both the parameter of variance and the parameter of truncation, which makes the analysis more convenient and the analytic formulas more concise. The truncation parameters for the inner bound and outer bound of $l^{2}$ norm of the codewords can be different, which will lead to the first and second order asymptotics with different coefficients. Nevertheless, it will not change the order of them as $O(n^{\frac{1}{2}})$ and $O(n^{\frac{1}{4}})$. In addition, the truncation parameter of the outer bound of radius will be reflected in the coefficient of these asymptotics as the power constraint is necessary in AWGN channels.
@@
Note that the truncation parameter $\mu$ lies in both the size of codebooks (6) and the size of the key (115), which is in contrast with the analysis of the first order asymptotics of covert communication over AWGN channels [10] [11] [12]. If we apply random coding with Gaussian distribution as in the achievability part of Theorem 5 in [12] or Theorem 9.1.1 in [38], we can not guarantee that the distribution of the codebook is Gaussian since we can not use $n\rightarrow\infty$ in the finite blocklength regime. The advantages of truncated Gaussian family as the generating distribution of the codewords are as follows. Firstly, all codewords satisfying power constraint is ensured so that we can control the power level easily. Secondly, the truncated Gaussian family is close to Gaussian family, and the effect of truncation on the output distribution is convenient to manipulate to satisfy the covert constraint.
Remark 14.
To satisfy the covert constraint with KL divergence in (12), the power level should be $\Psi(n)=O(\frac{1}{\sqrt{n}})$ in (113) of the condition (d). The kernel lies in the divergence between two Gaussian distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$. If the covert constraint is in the form of other divergences with proper upper bound $\delta$, such as total variational distance, Hellinger distance or $\chi^{2}$-divergence, the results will be almost the same except that $\Psi(n)$ may have a different constant coefficient. This fact will be illustrated later in Section V.
Remark 15.
For covert communication over AWGN channels, previous works [10] [11] have established the characterization of the first order asymptotics of the throughput as $O(\sqrt{n})$. This work focuses on refining these results by characterizing both the first and second order asymptotics as functions of the blocklength $n$, decoding error probability $\epsilon$, the background noise level parameter $\sigma$ and covert parameter $\delta$, which is not reported in [10] [11]. The analytic investigation stresses the influence of the ratio of the noise level (which is $\sigma^{2}/1$ in this work) on the throughput and the size of key bits.
Remark 16.
Our result on the size of key bits is slightly different from the case in discrete memoryless channels [22], and is also different from Theorem 6 in [10] where the numbers of key bits are $O(1)$ if the adversary is less noisy ($D_{Q}\leq D_{P}$). In this setting, the number of key bits also depends on the parameter of truncation $\mu$ with finite blocklength over AWGN channels. Even if the channel between Alice and Willie is severer than the main channel, the number of key bits should be of $O(\sqrt{n})$ if the noise level is only a bit larger. This is the necessary cost of power constraint for the sending codewords with finite blocklength over AWGN channels, which will lead to the truncation for Gaussian random coding in the finite blocklength regime.

### v6

@@ III-B Resolvability Analysis
$\begin{split}&\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{z},0})\\ =&\int_{z}\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{X}_{sw})\\ \times&\log\frac{\sum_{s=1}^{K}\sum_{w=1}^{M}W^{n}_{2}(\bm{z}|\bm{X}_{sw})}{MKQ^{(n)}_{\bm{Z},0}(\bm{z})}d\bm{z},\end{split}$ (70)
@@
Note that the three terms in (III-B) are all random variables as they are functions of $\bm{X}_{sw},s=1,\cdots,K,w=1,\cdots,M$, which are randomly generated from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$. Our strategy of bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ is as follows. We directly investigate the expectation of each term in (III-B). By showing that the expectation of each term in (III-B) is small, it is reasonable that the event that the realization of $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ induced by a set of legitimate codewords is small has positive probability. From the linearity of the expectation, we have
@@
Then, to upper bound the term (73), we divide the product space $\bm{X}\times\bm{Z}$ into two parts. For the first part, we show that the inner term $\log\left(1+\frac{W^{n}_{2}(\bm{Z}|\bm{X})}{MK\bar{Q}^{(n)}_{\bm{Z},1}(\bm{Z})}\right)$ for each $(\bm{x},\bm{z})$ pair is under control. For the second part, we prove that the probability of it is exponentially decreasing. In fact, a precise upper bounding of the first part depends on the choice of the sizes of codebook and the key, which will be illustrated later in Theorem 7.
Before establishing Lemma 9, we first define $\mathcal{E}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}\leq\gamma_{1}\}$. Note that in the definition of $\mathcal{E}$, we use ${Q}^{(n)}_{\bm{Z},1}$ but not $\bar{Q}^{(n)}_{\bm{Z},1}$. Then, some technical preparations are provided in Lemma 7 and Lemma 8. Lemma 7 is on the ratio $\frac{W_{2}^{n}(\bm{z}|\bm{x})}{{Q}^{(n)}_{\bm{Z},1}(\bm{z})}$, the proof of which can be found in Appendix D.
@@
For the event $\mathcal{E}^{c}=\{(\bm{x},\bm{z}):\log\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$, the following Lemma 8 provides an upper bound on its probability with proper chosen $\gamma_{1}$. Similar with the proof of Lemma 4, our strategy here is separately considering the codewords with equal power and choosing an universal threshold $\gamma_{1}$. As a consequence, the sphere symmetry can be leveraged to show that the probability of the event $\{\bm{z}:\frac{W_{2}^{n}(\bm{z}|\bm{x})}{Q^{(n)}_{\bm{Z},1}(\bm{z})}>\gamma_{1}\}$ for each $\bm{x}$ has large-deviation bounds.
@@
$\begin{split}&\frac{n}{2}\log\big(1+\frac{\mu\Psi(n)}{\sigma^{2}}\big)\\ +&\frac{\log e}{2}\sum_{i=1}^{n}\left[\frac{z_{i}^{2}}{\sigma^{2}+\mu\Psi(n)}-\frac{(z_{i}-\sqrt{\Upsilon})^{2}}{\sigma^{2}}\right]\\ =&\frac{n}{2}\log(1+\frac{\mu\Psi(n)}{\sigma^{2}})-\sum_{i=1}^{n}\frac{\log e}{2\sigma^{2}(\sigma^{2}+\mu\Psi(n))}\\ \times&\left[\mu\Psi(n)z_{i}^{2}-2(\sigma^{2}+\mu\Psi(n))\sqrt{\Upsilon}z_{i}+(\sigma^{2}+\mu\Psi(n))\Upsilon\right]\end{split}$ (81)
@@
The strategy of the proof of Lemma 8 is similar as Lemma 4. We first consider the codewords with the same $l^{2}$ norm $\sqrt{n\Upsilon}$ (on the sphere with the same radius or with equal power constraint), separately. For codewords on each sphere, we can utilize the sphere symmetry to simplify the analysis. Again, the parameter $\mu$ makes the interval under control so that an universal threshold $\gamma_{1}$ can be obtained. The inner probability $\mathbb{P}_{W_{2}^{n}(\bm{Z}|\bm{X}=\bm{x})}\bigg(\log\frac{W_{2}^{n}(\bm{Z}|{\bm{X}=\bm{x}})}{h_{1}^{(n)}(\bm{Z})}>\gamma_{1}\bigg)$ for each $\bm{x}$ is a tail probability and can be bounded by concentration inequality from Chernoff technique. As the bounds of the tail probability depends only on the radius of the codewords and are all exponentially decreasing with $n$, we can get the maximum of the tail probabilities for all $\bm{x}$, which is again exponential decreasing with $n$.
@@
In our previous work [23], the encoder adopts sequential random coding (Theorem 1 in [23]) so that the codewords are not i.i.d generated, which makes it difficult to perform resolvability analysis. In this work, the codewords are i.i.d sampled from the truncated Gaussian distribution $\bar{P}_{n}^{(n)}$ that is close to Gaussian family. The resolvability analysis on bounding the divergence between the induced output distribution of the codewords and the distribution of noise depends much on this fact.
Remark 10.
@@ III-C Existence of the Codebooks and Achievability Bound
Regarding the quantity $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ as a function of the codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$, we apply McDiarmid’s Inequality with Lemma 10 and obtain the following lemma. As the proof is simple, we omit it here.
@@
We have illustrated the random coding scheme, decoding method in Section III-A. The upper bound on the first type of decoding error in Lemma 4 and the average of the second type of decoding error has been analyzed in Lemma 6. We also bounded the expectation of the discriminative metric - $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. A legitimate covert code should be based on the previous Lemma 12. The choice of various parameters both reliability and covertness will be illustrated in our main results as next theorem to realize the event $\mathcal{E}_{1}\cap\mathcal{E}_{2}$. We will provide characterizations on the size $M_{n}$ of each codebook for covert communication over AWGN channels with sufficient large $n$ through proper choice of the power level $\Psi(n)$. The theorem also characterizes the necessary number of key bits $K_{n}$, and it is shown that $K_{n}$ depends not only on the noise level $\sigma^{2}$ at the adversary’s channel and covert parameter $\delta$, but also depends on the hyper-parameter of truncation $\mu$.
@@
Under the above conditions, with sufficiently large $n$, we can choose $c_{0}=13$, $c_{1}=1$, $c_{2}=11$, $c_{3}=4$, and $c_{5}=3$ and a proper power level $c_{4}=\frac{2\sigma^{2}}{\mu}\sqrt{\delta\ln 2}$. We obtain We obtain (1) and the number of key bits should satisfy (21).
@@
Based on Lemma 12, for each index $s$, we expurgate the $(1-\lambda_{2})M_{n}$ codewords with largest $\varepsilon_{sw}^{(2)}$ for from each sub-codebook $\{\bm{x}_{sw},w\in\langle M\rangle\}$. Thus,we get a new family of codebooks, each of which has size $\lambda_{2}M$. Without loss of generality, we assume that the remaining codewords have indexes from $1$ to $\lambda_{2}M$, and denote the maximum of error probability of the second type in these new codes as $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. These codewords with indexes from $\lambda_{2}M+1$ to $M$ have error probabilities of the second type at least larger than $\underset{w\in\langle\lambda_{2}M\rangle}{\max}\varepsilon_{sw}^{(2)}$. As the average error probability of the second type for each old codebook is less than $\lambda_{1}\frac{1}{n^{c_{0}}}\mathbb{E}_{\bar{Q}^{(n)}_{\bm{Y},1}}\bigg(\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}\bigg)$. We have for each $s$, the following inequality holds,
@@
from the conclusion of Lemma 4. For the covertness condition in (110), from (115) and Lemma 3, the term $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ satisfies
@@
$\log$ $\big(1+\frac{2^{\gamma_{1}+1}}{\lambda_{2}MK}\big)+An\cdot ae^{-bn}$ (121)
@@
With $\Psi(n)=c_{4}\cdot\frac{\sqrt{n}}{n}$, we get a lower bound of $\log M^{*}(n,\epsilon,\sigma,\delta)$ in (III-C).
$\log M^{*}(n,\epsilon,\delta)\geq$ $\frac{n}{2}\log(1+\mu c_{4}\frac{\sqrt{n}}{n})+\frac{(\mu^{2}-\mu)c_{4}\frac{\sqrt{n}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})}\cdot\log e+\sqrt{n\cdot\frac{2\mu^{2}c_{4}\frac{\sqrt{n}}{n}+\frac{c_{4}^{2}\mu^{2}}{n}}{2(1+\mu c_{4}\frac{\sqrt{n}}{n})^{2}}}$
$\times\Phi^{-1}(\epsilon_{1}-\frac{\max_{\Upsilon}B(\Upsilon,\mu\Psi(n))}{\sqrt{n}})\log e+O(\log n)$
@@
The condition in (113) is equivalent to
@@
$+$ $\sqrt{n\mathsf{V}(\mu^{2}\Psi(n),\mu\Psi(n))}\times\Phi^{-1}(1-\epsilon_{1})+c_{6}\log n\}$ (129)
@@
  • If $\sigma^{2}<\frac{1}{\mu^{2}}$, i.e., the channel between Alice and Willie is less noisy than the main channel, we have
@@
  • If $\sigma^{2}\geq\frac{1}{\mu^{2}}$, then $\phi(\mu,\sigma^{2})\leq 0$ and the first order asymptotics of the optimal number of key bits should be $O(1)$.
@@
If we apply random coding with Gaussian distribution as in the achievability part of Theorem 5 in [12] or Theorem 9.1.1 in [38], we can not guarantee that the distribution of the codebook is Gaussian since we can not use $n\rightarrow\infty$ in the finite blocklength regime. The advantages of truncated Gaussian family as the generating distribution of the codewords are as follows. Firstly, all codewords satisfying power constraint is ensured so that we can control the power level easily. Secondly, the truncated Gaussian family is close to Gaussian family, and the effect of truncation on the output distribution is convenient to manipulate to satisfy the covert constraint. The support set of the codewords is $\{\bm{x}:\sqrt{\mu^{2}n\Psi(n)}\leq\|\bm{x}\|\leq\sqrt{n\Psi(n)}\}$. The upper bound of $\|\bm{x}\|$ is for the power constraint, and we set the inner bound $\sqrt{\mu^{2}n\Psi(n)}$ here for upper bounding the error probability of the first type. If there exists codewords with too small $l^{2}$ norm, then there exists no $\gamma$ to upper bound the first type of error in decoding. In fact, the bounding of the first type of error in Lemma 4 relies on the fact that the power of each codeword lies in a finite interval (See Remark 5). Here we choose truncated Gaussian with decreasing variance $\mu\Psi(n)$. The parameter $\mu$ acts as both the parameter of variance and the parameter of truncation, which makes the analysis more convenient and the analytic formulas more concise. The truncation parameters for the inner bound and outer bound of $l^{2}$ norm of the codewords can be different, which will lead to the first and second order asymptotics with different coefficients. Nevertheless, it will not change the order of them as $O(n^{\frac{1}{2}})$ and $O(n^{\frac{1}{4}})$. In addition, the truncation parameter of the outer bound of radius will be reflected in the coefficient of these asymptotics as the power constraint is necessary in AWGN channels.
@@
To satisfy the covert constraint with KL divergence in (12), the power level should be $\Psi(n)=O(\frac{1}{\sqrt{n}})$ in (115) of the condition (d). The kernel lies in the divergence between two Gaussian distributions $Q^{(n)}_{\bm{Z},1}$ and $Q^{(n)}_{\bm{Z},0}$. If the covert constraint is in the form of other divergences with proper upper bound $\delta$, such as total variational distance, Hellinger distance or $\chi^{2}$-divergence, the results will be almost the same except that $\Psi(n)$ may have a different constant coefficient. This fact will be illustrated later in Section IV-C.
@@
Remark 15.
Our result on the size of key bits is slightly different from the case in discrete memoryless channels [22], and is also different from Theorem 6 in [10] where the numbers of key bits are $O(1)$ if the adversary is less noisy ($D_{Q}\leq D_{P}$). In this setting, the number of key bits also depends on the parameter of truncation $\mu$ with finite blocklength over AWGN channels. Even if the channel between Alice and Willie is severer than the main channel, the number of key bits should be of $O(\sqrt{n})$ if the noise level is only a bit larger. This is the necessary cost of power constraint for the sending codewords with finite blocklength over AWGN channels, which will lead to the truncation for Gaussian random coding in the finite blocklength regime. The truncation is not necessary when we only consider the first order asymptotics, such as the proof of the achievability part of Theorem 5 in [12]. This can be reflected by the following fact. If we let $\mu=\mu(n)=1-\frac{1}{n^{\tau}}$ with $0<\tau<\frac{1}{2}$, the behavior of $\Delta$ can be found as following Fig. 4. In this case, Corollary 1 still holds for sufficiently large $n$ and the other lemmas and theorems also hold. Under the circumstance, the codewords are far away from with equal power when $n$ is finite. The truncation effect $\mathbb{D}(\bar{P}_{n}^{n}\|P^{n}_{n})$ on the output distribution can be neglected in the expectation of the divergence $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ when $n$ is sufficiently large. Consequently, the first order and the second order asymptotics of the throughput are $\sigma^{2}(\delta n\log e)^{1/2}$ and $\sigma^{2}(\delta n\log e)^{1/2-\tau}$, respectively. In addition, the first- and second order asympotics of the number of key bits are $(1-\sigma^{2})\sqrt{n\delta\log e}$ and $O(n^{\frac{1}{2}-\tau})$, respectively. If we further let $\tau=\frac{1}{2}$, we can see that $\Delta$ is stationary with respect to $n$ from Fig .4 and can not be neglected. From the definition of $\bar{g}^{(n)}(\bm{x})$ and $g^{(n)}(\bm{x})$ and their relationship with $\bar{g}_{1}^{(n)}(\bm{y})$$g_{1}^{(n)}(\bm{y})$ and $\bar{h}_{1}^{(n)}(\bm{z})$$h^{(n)}_{1}(\bm{z})$, we can get that $\frac{\bar{g}_{1}^{(n)}(\bm{y})}{g_{1}^{(n)}(\bm{y})}$ and $\frac{\bar{h}_{1}^{(n)}(\bm{z})}{h^{(n)}_{1}(\bm{z})}$ are bounded for all $n$. From (68), the average of the error probability of the second type can be sufficiently small with same $M_{n}$. Meanwhile, Lemma 4 for the maximum of error probability of the first type still holds. Nevertheless, from the covertness condition (110), the covert condition holds only for $\{\delta:\delta>\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})=-\log\Delta\}$ since we can choose $\Psi(n)$ such that the other terms in (110) are sufficiently small for sufficiently large $n$. For these $\delta$, the second order asymptotics of $\log M^{*}(n,\epsilon,\sigma,\delta)$ is $O(n^{\frac{1}{4}})$. The additional cost on the covertness stems from the truncation. The above analysis implies that it is preferable to choose $\mu$ close to $1$ when the blocklength $n$ is large and a moderate $\mu$ between $\frac{1}{2}$ and $1$ when we have a moderate blocklength $n$. Nevertheless, as the blocklength $n$ is always finite, $\mu<1$ is necessary, which implies that the transmitter still needs keys to assist when $1<\sigma^{2}\leq\frac{1}{\mu^{2}}$.
Fig. 4: The value of $\Delta$ with proper chosen $\mu(n)=1-\frac{1}{n^{\tau}}$, which is from numerical computation of the analytic formula (32).
Our results reveal that the consumption of covertness comes from two aspects. The first is average power of the signal. From Theorem 5 of [12], Gaussian signals will result in least consumption of covertness. Nevertheless, we can not obtain perfect Gaussian signals in the finite blocklength regime. There is always some gap between the empirical distribution of signals and the Gaussian distribution. The second ingredient is that practical signals have finite power, which is the reason for truncation. These two aspects are reflected in two KL divergence terms in the inequality (110) as $\mathbb{D}(Q^{(n)}_{\bm{Z},1}\|Q^{(n)}_{\bm{Z},0})$ and $\mathbb{D}(\bar{P}_{n}^{(n)}\|P^{(n)}_{n})$, respectively. The key parameter of the two terms are $\mu$ and $\Psi(n)$. We will discuss on the choice of $\Psi(n)$ in Section IV later.